How many moles are in  1.29 x 1024 hydrogen atoms in HF?


1
Expert's answer
2021-02-01T04:04:04-0500

1 mol contains "6.022 \times 10^{23}" atoms.

Proportion:

1 mol – "6.022 \times 10^{23}"

x - "1.29 \times 10^{24}"

"x = \frac{1 \times 1.29 \times 10^{24}}{6.022 \times 10^{23}} \\

= 2.14 \;mol"

Answer: 2.14 mol